# Geometric Bounds for Persistence

### v1

@@ 2.3 Filling Radius
$\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))=\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))).$
This is not obvious by default, since the absolute filling radius in the right-hand side should be computed in the space $L^{\infty}(\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$. Nevertheless, this equality holds true. The inequality $\geq$” follows from the universal property. The inequality $\leq$” can be easily shown by making use of the evident $1$-Lipschitz restriction map $L^{\infty}(\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))\to L^{\infty}(\mathcal{X})$.
Auxiliary definitions. Before finishing the background section, we define two versions of the radius of a set $\mathcal{X}$, which will be used in the remainder of the paper. In the first one, there is no reference to an ambient space, and the center is in the set $\mathcal{X}$. In the second one (circumradius), the radius of $\mathcal{X}$ is computed in an ambient space, and the center may not be in $\mathcal{X}$.
@@
    $\mathrm{rad}(\mathcal{X})=\inf_{y\in\mathcal{X}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathcal{X}}(x,y).$
@@
    Assume additionally that $\mathcal{X}$ is a compact subset of a Banach space $\mathbf{Y}$, so that the metric $\mathbf{d}_{\mathcal{X}}$ agrees with the one inherited from $\mathbf{Y}$ (that is, the embedding $\mathcal{X}\hookrightarrow\mathbf{Y}$ is distance-preserving). The circumradius of $\mathcal{X}$ in $\mathbf{Y}$ is
    $\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})=\inf_{y\in\mathbf{Y}}\sup_{x\in\mathcal{X}}\|x-y\|.$
@@ 3 Bounding Lifespans via Widths
In this section, we recall and establish upper bounds the filling radius, which can be used to estimate the lifespans in both Čech and Vietoris–Rips settings. The basic idea behind many constructions is simple and can be illustrated as follows. Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$, and let $k>\ell$ be positive integers. A degree-$k$ homology class of $\mathcal{X}$ can be “killed” by deforming $\mathcal{X}$ to an $\ell$-dimensional complex inside $\mathbf{Y}$ (a “core”), and if every point moves by some controlled distance, then we get an estimate for the filling radius of the degree-$k$ homology.
@@ 3.1 Background on Widths
The Urysohn width was historically first. The following definition is equivalent to the one given by Pavel Urysohn around 1923 [Ale26] in the context of dimension theory.
@@
and $\operatorname{UW}_{0}(\mathcal{X})=\operatorname{diam}(\mathcal{X})$ if $\mathcal{X}$ is connected. The $n$-width of an $n$-dimensional manifold is zero, and all preceding $k$-widths are positive for $0\leq k<n$.
@@
The coefficients in the definition of $\rho(\mathcal{M})$ do not matter here; everywhere below we stick to the case of $\mathbb{Z}_{2}$-coefficients since this is the main case of interest for persistent homology. The importance of this result for bounding lifespans becomes immediate once one notices that the proof of this inequality can be easily generalized for any topological feature, and not just the fundamental class.
@@
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$,
@@
It is natural to look for a relative version of Theorem 3.3. We provide such an inequality below (Theorem 3.5), additionally replacing its right-hand side by the quantity that is comparable to the Urysohn width but easier to compute; the corresponding metric invariant implicitly appeared in the work of Pavel Alexandrov [Ale33] on dimension theory.
@@
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
@@
The Alexandrov width enjoys similar properties with the Urysohn width: it is monotonically decreasing in $k$ until it reaches zero when $k$ becomes equal the dimension of $\mathcal{X}$. If $\mathcal{X}$ is connected, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})$ equals the circumradius of $\mathcal{X}$ in $\mathbf{Y}$, that is, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})=\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ (see 2.35).
It is easy to see directly from the definitions that $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Combining this with Theorem 3.3, we get the estimate $\rho(\omega;\mathcal{X})\leq\operatorname{AW}_{k-1}(\mathcal{X}\subset\mathbf{Y})$, which we will improve by replacing the absolute filling radius by the relative one.
@@
This is the relative analogue of Theorem 3.3. It will follow from a stronger estimate below (Theorem 3.10), but the intuition behind it is simple as described at the preamble of this section: One can kill higher homology of $\mathcal{X}$ by deforming it to a low-dimensional complex in $\mathbf{Y}$, and if every point moves by some controlled distance, then we get an estimate for the filling radius.
The virtues of 3.5 are twofold. First, the Alexandrov width seems to be easier to estimate than the Urysohn width. Basically, the Urysohn width considers all maps to something $k$-dimensional, whereas the Alexandrov width only considers a special class of those maps with the images lying in the same ambient space $\mathbf{Y}$; this is also the reason why $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Second, in Appendix A we explain that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X})$. Hence, with 3.5, this implies one can bound the relative filling radius with the Urysohn width (if $\mathcal{X}$ is compact), too, and we do not lose much when substituting the widths, since
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. For an integer $k\geq 0$, the Kolmogorov $k$-width of $\mathcal{X}$ (relative to $\mathbf{Y}$) is defined as
$\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf\{r~{}|~{}\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{P}\subset\mathbf{Y})\text{ for some affine $k$-plane }\mathcal{P}\subset\mathbf{Y}\}.$
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Then for any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$, $k\geq 0$,
@@
$\mathrm{E}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{m}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1,\mbox{ and }x_{j}=0\mbox{ for }j>m\right\},$
@@
For more information on widths, we refer the reader to [Bal21].
@@ 3.2 Bounds for Čech Lifespans via Treewidth
where $\omega_{r}$ maps to $\omega$ by the homology map induced by the inclusion (see Section 2.3). To estimate lifespans, we need upper bounds for filling radii, such as in 3.5 and 3.7. Both estimates can be simultaneously strengthened using the following new width invariant. Recall that throughout the paper, we use homology with $\mathbb{Z}_{2}$-coefficients (though the same definition makes sense with any coefficients).
@@
$\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$, $k\geq 0$,
@@
Fix any number $\delta>\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$ and pick a witness map $f:\mathcal{X}\to\mathbf{Y}$ with at most $k$-dimensional image $\mathcal{T}$, such that $\|x-f(x)\|\leq\delta$ for all $x\in\mathcal{X}$. Pick a cycle $S\subset\mathcal{X}$ representing $\omega$ and continuously deform it via the linear homotopy $h:S\times[0,1]\to\mathbf{Y}$ given by
@@
Every point $x\in S$ moves by distance at most $\delta$, so the continuous deformation always stays in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\operatorname{\tilde{H}}_{k}(\mathcal{T})$. Therefore, $S$ is null-homologous in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
@@
Figure 1: A shape with small $1$-treewidth but considerable Alexandrov 00-width and Kolmogorov $1$-width
We proceed with a strengthening applicable to non-compact sets of the form $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Notice that the width in the right-hand side of the following estimate is computed on a compact set, since in Banach spaces the closure of the convex hull of a compact set is compact [Lax02, Section 13, Exercise 9].
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any non-negative number $r$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
@@
Let our homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is a compact subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (here $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity gives rise to a map from $U$ to the nerve $Z$ of that cover. By the nerve theorem this map $\phi:U\to Z$ is a homotopy equivalence, with an evident homotopy inverse $\psi:Z\to U$, which sends every vertex of $Z$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $Z$. Therefore, $\omega|_{U}=\psi_{*}\phi_{*}\omega|_{U}\in\operatorname{\tilde{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(Z)\subset\overline{\mathrm{conv}}(\mathcal{X})$. A similar argument shows that is $S$ bounds in $\mathcal{N}_{R}(\mathcal{X}\subset\mathbf{Y})$, for some $R>r$, then the filling chain can be also taken with the support in $\overline{\mathrm{conv}}(\mathcal{X})$. Therefore,
@@
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{TW}_{k}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y})$.
@@
Figure 2: For the surface $S$ above, while $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, the lifespan of red curve $\alpha\in\operatorname{\tilde{H}}_{1}(S)$ is large.
We would like to point out that while $k$-widths are effective for bounding lifespans of homology classes of degree $>k$, they do not say much about the lifespans of homology classes in lower degrees. In Figure 2, we give a simple $2$-dimensional example $S$. Here, $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, and hence, the homology class in $\operatorname{\tilde{H}}_{2}(S)$ has short lifespans by Section 1. However, one can easily see that homology classes $\alpha\in\operatorname{\tilde{H}}_{1}(S)$ can still have very large lifespans. One can generalize this example to any dimension and codimension. In Section 5 we take a different perspective, and give global bounds for lifespans and death times in all degrees simultaneously.

### v2

@@ 2.3 Filling Radius
$\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}=\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}.$
This is not obvious by default, since the absolute filling radius in the right-hand side should be computed in the space $L^{\infty}\big{(}\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\big{)}$. Nevertheless, this equality holds true, as it is explained in Section 3.3 (see 3.14).
Auxiliary definitions. Before concluding the background section, we define two versions of the radius of a set $\mathcal{X}$, which will be used in the remainder of the paper. In the first one, there is no reference to an ambient space, and the center is in the set $\mathcal{X}$. In the second one (circumradius), the radius of $\mathcal{X}$ is computed in an ambient space, and the center may not be in $\mathcal{X}$.
@@
    $\mathrm{rad}(\mathcal{X}):=\inf_{y\in\mathcal{X}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathcal{X}}(x,y).$
@@
    Assume additionally that $\mathcal{X}$ is a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. [^13] So that $\mathbf{d}_{\mathcal{X}}(x,x^{\prime})=\|x-y\|$ for all $x,x^{\prime}\in\mathcal{X}$ The circumradius of $\mathcal{X}$ in $\mathbf{Y}$ is
    $\mathrm{rad}(\mathcal{X}\subset\mathbf{Y}):=\inf_{y\in\mathbf{Y}}\sup_{x\in\mathcal{X}}\|x-y\|.$
@@ 3 Bounding Lifespans via Widths
In this section, we recall and establish upper bounds for the filling radius which can be used to estimate the lifespans in both Čech and Vietoris–Rips settings. The basic idea behind many constructions is simple and can be illustrated as follows. Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$, and let $k>\ell$ be positive integers. A degree-$k$ homology class of $\mathcal{X}$ can be “killed” by deforming $\mathcal{X}$ to an $\ell$-dimensional complex inside $\mathbf{Y}$ (a “core”), and if every point moves by some controlled distance, then we obtain an estimate for the filling radius of the degree-$k$ homology.
@@ 3.1 Background on Widths
The Urysohn width was historically the first one to be introduced. The following definition is equivalent to the one given by Urysohn around 1923 in the context of dimension theory; it was posthumously published by Alexandrov [Ale26].
@@
and $\operatorname{UW}_{0}(\mathcal{X}):=\operatorname{diam}(\mathcal{X})$ if $\mathcal{X}$ is connected. The $n$-width of an $n$-dimensional manifold is zero, and all preceding $k$-widths are positive for $0\leq k<n$.
@@
The homology coefficients in the definition of $\rho(\mathcal{M})$ do not matter here; everywhere below we assume them to lie in an arbitrary field, which is fixed and omitted from the notation. The importance of this result for bounding lifespans becomes immediate once one notices that the proof of this inequality can be easily generalized for any topological feature, and not just the fundamental class.
@@
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, and any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$,
@@
It is natural to look for a relative version of Theorem 3.3. We provide such an inequality below (Theorem 3.5), additionally replacing its right-hand side by the quantity that is comparable to the Urysohn width but easier to compute; the corresponding metric invariant implicitly appeared in the work of Alexandrov [Ale33] on dimension theory.
@@
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y}):=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
@@
The Alexandrov width enjoys similar properties with the Urysohn width: it is monotonically decreasing in $k$ until it reaches zero when $k$ becomes equal the dimension of $\mathcal{X}$. If $\mathcal{X}$ is connected, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})$ equals the circumradius of $\mathcal{X}$ in $\mathbf{Y}$, that is, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})=\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ (see 2.39).
It is easy to see directly from the definitions that $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Combining this with Theorem 3.3, we obtain the estimate $\rho(\omega;\mathcal{X})\leq\operatorname{AW}_{k-1}(\mathcal{X}\subset\mathbf{Y})$, which we will improve by replacing the absolute filling radius by the relative one.
@@
This is the relative counterpart of Theorem 3.3. It will follow from a stronger estimate below (Theorem 3.10), but the intuition behind it is simple as described at the preamble of this section: One can kill higher homology of $\mathcal{X}$ by deforming it to a low-dimensional complex in $\mathbf{Y}$, and if every point moves by some controlled distance, then we obtain an estimate for the filling radius.
The virtues of 3.5 are twofold. First, the Alexandrov width seems to be easier to estimate than the Urysohn width. Basically, the Urysohn width considers all maps to a $k$-dimensional space, whereas the Alexandrov width only considers a special class of those maps with the images lying in the same ambient space $\mathbf{Y}$; this is also the reason why $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Second, in Appendix A we explain that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X})$. Hence, together with 3.5, this implies one can bound the relative filling radius with the Urysohn width (if $\mathcal{X}$ is compact), too, and we do not lose much when substituting the widths, since
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. For an integer $k\geq 0$, the Kolmogorov $k$-width of $\mathcal{X}$ (relative to $\mathbf{Y}$) is defined as
$\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}):=\inf\{r~{}|~{}\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{P}\subset\mathbf{Y})\text{ for some affine $k$-plane }\mathcal{P}\subset\mathbf{Y}\}.$
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Then for any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$, $k\geq 0$,
@@
$\mathrm{E}:=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{m}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1,\mbox{ and }x_{j}=0\mbox{ for }j>m\right\},$
@@
For more information on widths, we refer the reader to Balitskiy [Bal21].
@@ 3.2 Bounds for Čech Lifespans via Treewidth
where $\omega_{r}$ maps to $\omega$ by the homology map induced by the inclusion (see Section 2.3). To estimate lifespans, we need upper bounds for filling radii, such as in 3.5 and 3.7. Both estimates can be simultaneously strengthened using the following new width invariant. Recall that throughout the paper, we use homology with coefficients in a fixed field, which is omitted from the notation.
@@
$\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y}):=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, and any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$,
@@
Fix any number $\delta>\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$, and pick a witness map $f:\mathcal{X}\to\mathbf{Y}$, whose image $f(\mathcal{X})=\mathcal{T}$ has dimension at most $k$ and trivial $\widetilde{\mathrm{H}}_{k}(\mathcal{T})$, such that $\|x-f(x)\|\leq\delta$ for all $x\in\mathcal{X}$. Pick a cycle $S\subset\mathcal{X}$ representing $\omega$ and continuously deform it via the linear homotopy $h:S\times[0,1]\to\mathbf{Y}$ given by
@@
Every point $x\in S$ moves by distance at most $\delta$, so the continuous deformation always stays in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\widetilde{\mathrm{H}}_{k}(\mathcal{T})$. Therefore, $S$ is null-homologous in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
@@
Figure 1: A shape with small $1$-treewidth but considerable Alexandrov 00-width and Kolmogorov $1$-width.
We proceed with a strengthening applicable to non-compact sets of the form $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Notice that the width in the right-hand side of the following estimate is computed on a compact set, since in Banach spaces the closure of the convex hull of a compact set is compact; see e.g. Lax [Lax02, Section 13, Exercise 9].
@@
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any positive number $r$, and any homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
@@
Let our homology class $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is a compact subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (recall that $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity gives rise to a map from $U$ to the nerve $L$ of that cover. By the nerve theorem, this map $\phi:U\to L$ is a homotopy equivalence, with an evident homotopy inverse $\psi:L\to U$, which sends every vertex of $L$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $L$. Therefore, $\omega|_{U}=\psi_{*}\circ\phi_{*}(\omega|_{U})\in\widetilde{\mathrm{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(L)\subset\overline{\mathrm{conv}}(\mathcal{X})$. A similar argument shows that if $S$ bounds in $\mathcal{N}_{R}(\mathcal{X}\subset\mathbf{Y})$, for some $R>r$, then the filling chain can also be taken with the support in $\overline{\mathrm{conv}}(\mathcal{X})$. Therefore,
@@
In particular,
$d_{\omega}-b_{\omega}\leq\operatorname{TW}_{k}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y}).$
@@
Figure 2: For the surface $S$ above, while $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, the lifespan of red curve $\alpha\in\widetilde{\mathrm{H}}_{1}(S)$ is large.
We would like to point out that while $k$-widths are effective for bounding lifespans of homology classes of degree $>k$, they do not say much about the lifespans of homology classes in lower degrees. In Figure 2, we give a simple $2$-dimensional example $S$. Here, $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, and hence, the homology class in $\widetilde{\mathrm{H}}_{2}(S)$ has short lifespans by Section 1.2. However, one can easily see that homology classes $\alpha\in\widetilde{\mathrm{H}}_{1}(S)$ can still have very large lifespans. One can generalize this example to any dimension and codimension. In Section 5 we take a different perspective, and give global bounds for lifespans and death times in all degrees simultaneously.